OpenStudy (anonymous):

Does the function satisfy the hypotheses of the Mean Value Theorem on the given interval? f(x) = 2x^2 − 3x + 1, [0, 2] I found that c=3/4 but it was marked as wrong. Can you help?

OpenStudy (anonymous):

it has to satisfy it because it is a polynomial and therefore continuous and differentiable everywhere

OpenStudy (anonymous):

because it is a quadratic polynomial (degree 2) the number that satisfies the hypothesis of the MVT i.e. the \(c\) with \(f'(c)=\frac{f(b)-f(a)}{b-a}\) is actually the center of the interval

OpenStudy (anonymous):

we can go through the work to find it if you like, but odds on it is 1

OpenStudy (anonymous):

did you find \(f(2)\) ?

OpenStudy (anonymous):

yes, f(2) =3

OpenStudy (anonymous):

ok and \(f(0)=1\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \(f(2)-f(0)=3-1=2\)

OpenStudy (anonymous):

and therefore \(\frac{f(b)-f(a)}{b-a}=\frac{3-1}{2-0}=\frac{2}{2}=1\)

OpenStudy (anonymous):

\[f'(x)=4x-3\] set \[4x-3=1\]solve for \(x\) and i bet you get \(x=1\)

OpenStudy (anonymous):

awesome. Thanks!

OpenStudy (anonymous):

yw any time you have a quadratic, the number you are looking for is always the center of the interval you can prove it for yourself by considering a general quadratic \[f(x)=ax^2+bx+c\]